ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 224, 2018

Aurichi, L. F., A. Bella and R. R. Dias Tightness games with bounded finite selections

133

Bella, A. See L. F. Aurichi Belliart, M. A differentiable classification of certain locally free

actions of Lie groups

315

Ben Yaacov, Ï. On a Roelcke-precompact Polish group that cannot act transitively on a complete metric space

105

Berkowitz, R. and P. Devlin A stability result using the matrix norm to bound the permanent

437

Blair, M. D. On logarithmic improvements of critical geodesic restriction bounds in the presence of nonpositive curvature

407

Bodirsky, M., D. Evans, M. Kompatscher and M. Pinsker A counterexample to the reconstruction of ω-categorical structures from their endomorphism monoid

57

Bonatti, C. and R. Potrie Many intermingled basins in dimension 3

293

Cody, B. and M. Eskew Rigid ideals  343

Dahinden, L. Lower complexity bounds for positive contactomorphisms

367

Devlin, P. See R. Berkowitz Dias, R. R. See L. F. Aurichi Eldan, R., F. Nazarov and Y. Peres How many matrices can be

spectrally balanced simultaneously?

385

Eskew, M. See B. Cody Evans, D. See M. Bodirsky Hieronymi, P. and E. Walsberg Interpreting the monadic second

order theory of one successor in expansions of the real line

39

Janvresse. É, E. Roy and T. de la Rue Invariant measures for Cartesian powers of Chacon infinite transformation

1

Kompatscher, M. See M. Bodirsky Koszmider, P. Uncountable equilateral sets in Banach spaces of the

form C(K) 83

Lemmens, B., B. Lins and R. Nussbaum Detecting fixed points of nonexpansive maps by illuminating the unit ball

231

Lins, B. See B. Lemmens Malliaris, M. and S. Shelah Keisler’s order has infinitely many

classes

189

Nazarov, F. See R. Eldan Nuño-Ballesteros, J. J., B. Oréfice-Okamoto and J. N.

Tomazella Erratum to “The vanishing Euler characteristic of an isolated determinantal singularity”, Israel J. Math. 197 (2013), 475–495

505

Nussbaum, R. See B. Lemmens Nystedt, P., J. Öinert and J. Richter Non-associative Ore

extensions

263

Öinert, J. See P. Nystedt Oréfice-Okamoto, B. See J. J. Nuño-Ballesteros Peres, Y. See R. Eldan Pinsker, M. See M. Bodirsky Potrie, R. See C. Bonatti Richter, J. See P. Nystedt Roy, E. See É. Janvresse de la Rue, T. See É. Janvresse Shelah, S. See M. Malliaris

Shen, Z. and G. Yang On Square metrics of scalar flag curvature  159

Temkin, M. Functorial desingularization over Q: boundaries and the embedded case

455

Tomazella, J. N. See J. J. Nuño-Ballesteros Walsberg, E. See P. Hieronymi Yang, G. See Z. Shen